The equity sections from Atticus Groups 2016 and 2017 year-end balance sheets follow. Stockholders Equity (December 31,2016) Common stock-$4 par valuo, 100,000 shares Paid-in capital in excess of par value, common stock Retained eanings Total stockholders equity 120,000 $600,000 Stockholders Equity (December 31, 2017) Common stock-$4 par value, 100,000 shares $189,600 authorized, 47400 shares issued, 3,000 shares in treasury Paid-in capital in axcoss of par value, common stock Rotained oamnings ($30,000 restrictod by troasury stock 400,000 68.800 300001 $738,800 Less cost of treasury stock Total stockholdors equity The following transactions and events affected its equity during year 2017 Jan. 5 Mar. 20 Apr. 5 July 5 July 31 Aug. 14 Oct. 5 Declared a S0.50 per share cash dividend, date of record January 10. Purchased treasury stock for cash. Declared a S0.50 per share cash dividend, date of record April 10. Declared a $0.50 per share cash dividend, date of record July 10. Declared a 20% stock dividend when the stocks market value was $12 per share. Issued the stock dividend that was declared on July 31 Declared a S0.50 per share cash dividend, date of record October 10. Required 1. How many common shares are outstanding on each cash dividend date? 2. What is the total dollar amount for each of the four cash dividends? 3. What is the amount of the capitalization of retained earnings for the stock dividend? 4. What is the per share cost of the treasury stock purchased? 5. How much net income did the company earn during year 2017?
Answer #1

Solution 1:

Common share outstanding on each cash dividend date:

Jan 5 : 40000 shares

Apr 5 = Share issued - Shares in treasury stock = 40000 - 3000 = 37000 shares

July 5 = Share issued - Shares in treasury stock = 40000 - 3000 = 37000 shares

Oct 5 = Share issued - Shares in treasury stock = 47400 - 3000 = 44400 shares

Solution 2:

Computation of dollar amount for each cash dividend:

Jan 5 : Outstanding shares * Dividend per share = 40000 * $0.50 = $20,000

Apr 5 = 37000*$0.50 = $18,500

July 5 = 37000*$0.50 = $18,500

Oct 5 = 44400*$0.50 = $22,200

Solution 3:

Amount of capitalization of retained earnings for the stock dividend = Nos of shares issued in stock dividend * Market price per share on date of declaration = (37000*20%) * $12 = $88,800

Solution 4:

Per share cost of treasury stock purchased = Total cost of treasury stock / Nos of share in treasury stock

= $30,000 / 3000 = $10 per share

Solution 5:

Retained earning 2017 = Retained earnings in 2016 + Net Income for 2017 - Cash dividend paid - Stock dividend

$400,000 = $320,000 + Net Income for 2017 - ($20,000 + $18,500 + $18,500 + $22,200) - $88,800

Net Income for 2017 = $248,000